2.0mol of (NH4)3PO4
calculate the grams of phosphate (PO4) ions
1
Expert's answer
2020-04-27T01:44:58-0400

Solution.

"\nu((NH_4)_3PO_4) = 2.0 mol;"

"m = \nu\sdot M;"

"M(PO_4^3- ) = 31 + 16\sdot4 = 95 g/mol;"

"2.0 mol (NH_4)_3PO_4 contains 2.0mol PO_4^3-;"

"m(PO_4^3-) = 95 g/mol \sdot 2.0 mol = 190g;"

Answer: "m(PO_4^3-) = 190g."